OpenStudy (anonymous):

use the method of sucessive differences to determine the next number in the given sequence -3, -3, 3, 15, 33, 57, 87

OpenStudy (paxpolaris):

-3-(-3)=0 3 - (-3)=6 15-3 =12 33-15=18 ...